OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = 6x2 – 24x + 32?

OpenStudy (anonymous):

(–2, –8) (–2, 8) (2, –8) (2, 8)

OpenStudy (kainui):

Take the derivative by the power rule which is just taking the exponent and multiplying the coefficient by it, then subtracting the exponent by 1 to get, y'=12x-24. Since you know the slope is 0 at the vertex, plug that into y' and solve for x, which is 2. Plug 2 back into your original function and find y, simple as that.

OpenStudy (anonymous):

so its C or D

OpenStudy (anonymous):

?? I'm still confused

OpenStudy (anonymous):

i don't know which is the correct answer

OpenStudy (anonymous):

ohh wait is it (2,8)?

OpenStudy (anonymous):

-2,8

OpenStudy (anonymous):

you sure? y=6(x^2-4x)+32 y=6(x-2)^2-24+32 y=6(x-2)^2+8 so, the vertex is (2,8) is B!!!!!

OpenStudy (anonymous):

I'm 99% it's -2 but if you're confident with you answer, go with it! :)

OpenStudy (anonymous):

*99% sure

OpenStudy (anonymous):

ok thanks ! :)